OpenStudy (anonymous):

help simplify (m^2)(m^7)(2n)(5n)

OpenStudy (bahrom7893):

(m^2) * (m^7) * (2n) * (5n) = [m^(2+7)] * 10n^2 = 10 * n^2 * m^5

OpenStudy (anonymous):

which is 10m^9nm is it not? or is it simply left at (m^9)(10n^2)?

OpenStudy (bahrom7893):

oh yeah 9, sorry...

OpenStudy (bahrom7893):

good job math freak

OpenStudy (bahrom7893):

2+7 is nine, lol not 5!

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

no prob, lol

OpenStudy (bahrom7893):

lol man i just got back from college calculus, brains still flying around

OpenStudy (anonymous):

no prob, happens to me too, but with precal, haha